Prev Next

SW&IT Meeting Minutes 2022-05-09

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20220502 - Bec, Matthieu 80.0
LR20220502 - Mechtley, Brandon 72.0
LR20220502 - Eychaner, Glenn 32.0
LR20220502 - Edmeades, David 40.0
LR20220502 - Grenz, Paul 12.0
LR20220502 - Hooper, Stephen 0.0
LR20220502 - Zhang, Yang 40.0
LR20220502 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • Regarding last week's discussion of obs computers: we might want to have an obs computer in Steward, because the LUCI GUIs should perform better when run remotely from a computer in Steward than a computer on the mountain.

Team reports


  • Last week
    • OC, Management, AOSWSYS, ATT, Luci-gui-WG
    • GMT M1 report
    • SOUL comm plan and documents from EPinna review
    • AOS RTDB disconnect diagnostic patch
    • Misc AoArcetriQA entries
  • This week
    • ARGOS github synch (Lorenzo?)
    • MG support contract amends
    • AzCam kickoff with M. Lesser
    • LN SOUL compat, meet Florian?
    • Investigate RunAO timeout propagation (IT8573) ctd.


  • Previous:
    • GCS patch and AOS patch (with Brandon)
    • TMS software work and data analysis and debugging.
    • LBC flat program and wiki issue investigation.
    • Some issuetrak investigation.
    • Some safety training.
  • Next:
    • TMS wiki work and algorithm discussions with Andrew and John.
    • TMS software work.
    • Some LBC/MODS issues to follow-up with Olga.
    • LBC code study.
    • Several issuetrak investigation.
    • Investigate adding new image support into DMS Python APIs.
    • More safety training.


  • Last week
    • Finished up backup server fix, back to functionality
    • Got a sample LBTI user imported to LBTO GWorkspace
      • Massive pain, typical Google crap, glad we only have 20 users
    • IssueTrak reports locating a bug triggered by our critical ticket notifications, has a fix in the pipe
    • New IT API; will investigate it and see if there are better tools for the critical notifications
  • This week
    • Add reporting for RAID health on backup server
    • Keep working on LBTI
    • AC Vault access; last I heard he needed a PW reset but he stopped communicating. I'll ping him and see where he is.


  • Last week
    • Updated documentation for ARGOS.
    • Several Bug fixes for SHARK-NIR
    • LMIRCAM Bug fixes
    • Phasecam testing and bug fixes
  • This week
    • AZCAM talk with Mike Lesser
    • Continuing SHARK-NIR fixes
    • Continuing phasecam testing


  • Last week:
    • Updated LUCI Summer Shutdown 2022 plans.
    • Closed IT 8619 (problem centering slit).
    • Created a test case in my simulated TCS for IT 8573. Results added to IT notes.
    • First meeting of the LUCI GUI Working Group. Updated the Plan of Intent.
    • Sent my notes on TigerVNC to the LUCI GUI Working Group and IT.
    • Received info from R.Mather indicating that IT 8603 might be an electronics issue. Next action set to D.Thompson.
    • Worked half days Monday and Tuesday while my oldest was in town.
  • This week:
    • Fix the LUCI bugs revealed by the testing of IT 8573.
    • Work on Observer Scheduler GUI (IT 8571, 8529).
    • Second meeting of the LUCI GUI Working Group Wednesday at 4pm.


  • Last week
    • study MMC actuator list
    • support LBTI DX daytime work
    • Switch DX ASM configuration for day and night
    • support IF and disturb files generation
    • check/study new KL of DX ASM base files
    • ao_stop process improvement
  • This week
    • study and create the MMC actuator list
    • improveand install new KL of DX ASM base files
    • Switch DX ASM configuration for day and night
    • support LBTI DX daytime work (also first night)


  • Last week
    • vm testing: setncpa, gain ramp
    • ao-supervisor release tested on sxlbti, sxadsec, sxsoul--need to upgrade stable.
    • JCG saw more ice dispatch errors--should make an IT maybe.
    • support saturday/sunday: none needed, noticed some problems on WFS boot with JCG (TTCtrl & Ocam). He'll need to make an IT.
  • This week
    • set_z0 test?
    • setncpa save fits file, will PR to alfio's ncpa branch
    • VM gain ramp testing (have it executing the IDL--added a line so it doesn't execute write/read_seq in simulation--no mirrorctrl to read from/write to anyway, traced rip/disablecoils, adam communication won't do anything in simulation)
    • look at soul commissioning report from enrico
    • more issues:
      • aos rtdb disconnects, dropped rtdb messages when loading disturbances
      • wfsarb: on acquirerefao, set CCD viewer overly message only after baysides/filterwheel have moved
      • flaogui: disable "interrupt" button during uinterruptible commands
      • adsecarb: thread safety issue


  • Fedora 35 docker and system's RPM builds
  • Move room 530 to room 507
  • Problems with 507-3 (couldn't figure it out)
  • OVMS telemetry restart
  • PEPSI network configuration
  • UITS network stuff


  • core switches
  • Zabbix server
Topic revision: r11 - 09 May 2022, BrandonMechtley
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback